OpenStudy (anonymous):

One of the roots of the equation x2 + 3x - 18 = 0 is 3. What is the other root?

OpenStudy (anonymous):

Is that supposed to be x^2?

OpenStudy (anonymous):

x^2+3x-18=0 (x+6)(x-3)=0 so x=3 or x=-6

OpenStudy (anonymous):

Oh my god, I need to go drink more coffee!

OpenStudy (anonymous):

(x-3)(x+6) x multiplied by x= x^2 x multiplied by 6 = 6x -3 multiplied x = -3x and lastly -3 multiplied by 6= -18 so you can add-3x and 6x to get 3x and you'll be left with x^2+3x-18=0